OpenStudy (anonymous):

An ordinary (fair) die is a cube with the numbers through on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment. Compute the probability of each of the following events: Event : The sum is greater than 5 . Event : The sum is divisible by 5.

OpenStudy (anonymous):

Total possible outcomes are 36 Out of these 26 will give sum greater than 5 so P(sum greater than 5) = 26/36 = 13/18

OpenStudy (anonymous):

For 7 outcomes the sum is divisible by 5 So P(sum is divisible by 5) = 7/36 Hope it is clear.....???